Circular No. 8528

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 J2 (CATALINA) An apparently asteroidal object discovered by the Catalina Sky Survey

(discovery position below), and placed on the ‘NEO Confirmation Page’, was reported by J. Young (Table Mountain Observatory, 0.6-m reflector + CCD) on May 13.3 UT as being 3′′ across and possibly cometary, although this was far from certain because of extensive cirrus. Marginal-quality 60second exposures at 1.7 airmasses by E. Beshore (Mount Lemmon Survey, 1.5-m reflector + CCD) on May 14.2 UT showed a possible faint coma 5′′–7′′ across and a faint tail about 10′′ long in p.a. ∼ 100^{o}.

2005 UT  α_{2000}  δ_{2000}  Mag. Observer
May 12.16068 11^{h}35^{m}53.^{s}42 −13^{o}12′58′′.1 18.9 Catalina

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris are given on MPEC 2005-K01.

T = 2005 Mar. 30.743 TT  ω = 198.^{o}905
Ω = 33.360 }2000.0
q = 4.28168 AU  i = 150.790

COMET P/2005 JN (SPACEWATCH) An apparently asteroidal object reported by the Spacewatch program

(discovery observation below, cf. MPS 133394) has been found to show a coma with diameter 6′′ and a 20′′ tail in p.a. 145^{o} on R-band CCD frames taken by C. W. Hergenrother (Lunar and Planetary Laboratory) with the University of Arizona 1.54-m Kuiper reflector at Catalina on May 12.1 UT.

2005 UT  α_{2000}  δ_{2000}  Mag.
May 3.20314  13^{h}22^{m}55.^{s}34  +2^{o}13′14′′.9  20.4

The most recent astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2005-K11.

T = 2005 June 20.7004 TT  ω = 153.^{o}7724
e = 0.349979  Ω = 70.8036 }2000.0
q = 2.274666 AU  i = 8.8549
a = 3.499375 AU n^{o} = 0.1505630 P = 6.55 years
2005 May 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green